OpenStudy (anonymous):

(y/x)+(x/y)=2y calculate the derivative respect to x (implicit differentiation)

OpenStudy (anonymous):

\[\frac{ y }{ x } + \frac{ x }{ y } = 2y\] \[\frac{ y^2 + x^2 }{ xy } = 2y\] \[y^2 + x^2 = (xy) *2y\] \[y^2 + x^2 = 2xy^2\]